# Why do zeroes happen? A model-based approach for demand classification

### v1

@@ Abstract
Effective demand forecasting is critical for inventory management, production planning, and decision making across industries. Selecting the appropriate model and suitable features to efficiently capture patterns in the data is one of the main challenges in demand forecasting. In reality, this becomes even more complicated when the recorded sales have zeroes, which can happen naturally or due to some anomalies, such as stockouts and recording errors. Mistreating the zeroes can lead to the application of inappropriate forecasting methods, and thus leading to poor decision making. Furthermore, the demand itself can have different fundamental characteristics, and being able to distinguish one type from another might bring substantial benefits in terms of accuracy and thus decision making. We propose a two-stage model-based classification framework that in the first step, identifies artificially occurring zeroes, and then classifies demand to one of the possible types: regular/intermittent, intermittent smooth/lumpy, fractional/count. The framework utilises statistical modelling and information criteria to detect anomalous zeroes and then classify demand into those categories. We then argue that different types of demand need different features, and show empirically that they tend to increase the accuracy of the forecasting methods compared to those applied directly to the dataset without the generated features and the two-stage framework. Our general practical recommendation based on that is to use the mixture approach for intermittent demand, capturing the demand sizes and demand probability separately, as it seems to improve the accuracy of different forecasting approaches.
@@ 1 Introduction
People working in the area of demand forecasting sometimes encounter zeroes in their data. These zeroes can happen for a variety of reasons: some of them occur naturally, because no one buys the product in that specific moment of time, some happen artificially due to problems in inventory management system or recording errors. It is important to distinguish these two situations and treat them differently. In case of the naturally occurring zeros, we typically have intermittent demand, which according to the definition of Svetunkov and Boylan (2023) is the demand that “has non-zero values occurring at irregular frequency”. Many statistical methods have been developed, starting from the Croston (1972) and its bias corrected form by Syntetos and Boylan (2001) to machine learning methods developed over the years (Hasni et al., 2018; Babai et al., 2020; Jiang et al., 2020), so an analyst can select the most appropriate or the favourite approach and use it for intermittent demand forecasting. However, if we deal with the artificially occurring zeroes, they need to be treated differently: for example, using an intermittent demand approach on the data with stockouts would be harmful for decision making, because we would be forecasting stockouts instead of demand.
@@
    suggesting several fundamental features based on (1) and (2) that, as we argue, should improve the performance of forecasting approaches.
@@ 2.1 What is “intermittent demand”?
The rise of the interest in the area of intermittent demand started with the paper of Croston (1972), who acknowledged that the simple exponential smoothing (SES by Brown, 1956) is not appropriate when used on the data with unpredictable zeroes. To solve the problem, he suggested to split the data into two time series: demand sizes and demand intervals. His idea was that the demand that we observe ($y_{t}$) can be represented as a combination of two variables:
@@
where $z_{t}$ is the demand size on observation $t$ and $o_{t}$ is a binary variable of demand occurrence, that has a probability of occurrence $p_{t}=\frac{1}{q_{t}}$, where $q_{t}$ is the interval between the observed demand sizes. To produce forecasts, Croston (1972) used two simple exponential smoothing methods for capturing dynamics of the demand sizes and demand intervals. While being efficient and innovative, his approach was neglected by academia for more than 20 years until Willemain et al. (1994) and Johnston and Boylan (1996) showed that Croston’s method performed well in practice and should be preferred for intermittent demand instead of other simple forecasting methods. Acknowledging the existence of intermittent demand, these papers also opened a new direction of research – intermittent demand forecasting, where the patterns of the data are so different that the conventional forecasting techniques might fail or not work efficiently.
@@
From the practical point of view, when separating the regular demand from the intermittent, one still faces a challenge, because there are no appropriate rules and it is not clear, what quantity of zeroes transforms the regular demand into intermittent. Some practitioners use arbitrary thresholds of 10%, 15%, 20% etc of zeroes in the data as cut-off points, where one would need to switch from regular to intermittent demand forecasting method. But those thresholds do not have any theoretical rationale behind them, and can only be considered as approximations to the real solution of the problem. Even non-OR methods that could be used to forecast special events are still based on some arbitrary threshold values which differ across domains (Nikolopoulos, 2021). So far, the assumed “rule of thumb” is that any quantity of zeroes implies that the demand is intermittent.
@@
Furthermore, the academic definitions of intermittent demand found in the literature are generally broad. Early definitions, such as Silver (1981), describe it as demand with “significant periods of no demand activity” but the term “significant” is unclear. Later definitions, including Willemain et al. (1994) and Syntetos and Boylan (2001), describe it as random demand with many (or large proportion of) zero periods, but the terms “many” and “large proportion” used remain ambiguous. More recent definitions of Syntetos and Boylan (2005), Syntetos et al. (2009), Teunter et al. (2011), and Babai et al. (2014) describe intermittent demand as appearing sporadically, which is, in essence, correct, but does not distinguish between random absence of demand and systematic gaps caused by external factors such as disruptions or recording errors.
@@
Lately, some definitions became more case-specific and sometimes confusing, like “the features of intermittent demand are characterised by their irregularity, with a very small demand size” (Jiang et al., 2020), or that it is “characterised by time series with many zeroes” (Prestwich et al., 2021), or characterised by “irregular demand occurrence and low demand quantity variation” (Rožanec et al., 2022).
@@
One other mistake sometimes made in the literature is equating count demand to the intermittent one (Snyder et al., 2012). While in some situations this is correct, this is not universally the case because the term “intermittent” relates to the demand intervals, while the word “count” describes demand sizes. In reality, there are many examples of demand being intermittent yet fractional (e.g. electricity vehicles charging).
@@
Among all definitions, we find the one by Svetunkov and Boylan (2023) to be the most concise and clear: “intermittent time series is a series that has non-zero values occurring at irregular frequency”. This definition shows that the non-zero values happen randomly and cannot be predicted, and the definition does not impose any subjective terms like “many”, or “large”, or “some”. However, it still does not make a distinction between naturally occurring zeroes and the ones happening due to external factors.
@@ 2.2 Demand classification
There are several proposed classification schemes based either on variance partition, or the accuracy of forecasting procedures (Babiloni et al., 2010). We discuss the main ones here.
Williams (1984) was one of the first who proposed classifying demand into ‘smooth’, ‘slow-moving’ and ‘sporadic’ by partitioning the variance of demand over the lead time into causal parts. Eaves and Kingsman (2004) expanded this classification by inclusion of the irregular type, which is differentiated from the smooth one according to the level of demand sizes variability. However, both papers clearly state that the boundaries between these categories are industry/data-specific, and they should come as a managerial decision.
Syntetos et al. (2005) proposed a classification (now called ‘SBC’, Syntetos-Boylan-Croston) based on other principles. They discussed the existing categorisation schemes and compared Croston, Simple Exponential Smoothing (SES) and Syntetos-Boylan Approximation (SBA from Syntetos and Boylan, 2005) based on their theoretical Mean Square Error (MSE) values. The authors showed that one can derive the cut-off values for average demand intervals and coefficient of variation, allowing to select between Croston and SBA for each type of demand. This scheme has gained large popularity among practitioners, because it suggest four distinct categories of intermittent demand:
@@
In many cases practitioners use the scheme as a prior step for data analysis, categorising the intermittent demand into these categories without any specific purpose, completely neglecting it in the following steps, where, for example, they use some machine learning methods for demand forecasting. We should point out that the original motivation of Syntetos et al. (2005) was to help in choosing between Croston and SBA, rather than doing a categorisation for the sake of it. This idea seems to have been lost over the years.
The modification of the SBC was proposed by Kostenko and Hyndman (2006), who modified it, using an inequality from the ratio of MSEs from the original paper of Syntetos et al. (2005). They demonstrated that the parameter space for different forecasting methods has a non-linear cut-off. This classification, while being mathematically correct, has not gained as much popularity as SBC, being more complex.
Finally, Petropoulos and Kourentzes (2015) proposed a further refinement of the Kostenko and Hyndman (2006) scheme by adding SES method to the classification. They showed on the example of the Royal Airforce data (RAF) that this addition produces more accurate and less biased forecasts than the Kostenko and Hyndman (2006) scheme.
However, all the research that has been done in this direction up to this point has mainly focused on selecting between Croston and SBA specifically for intermittent demand. This implies that if there is at least one zero in the data, then the data can be flagged as intermittent and either of these two methods should be used. While this is widely true, there is evidence from Syntetos and Boylan (2006) that SES and Simple Moving Average (SMA) perform well even on data with some zeroes. So a refined approach is needed in order to classify demand as intermittent or regular.
@@ 2.3.1 Exponential smoothing
Research on exponential smoothing for intermittent demand has advanced from early explorations of Croston’s method and its statistical foundations (Shenstone and Hyndman, 2005) to more refined models tailored to count and intermittent data. Hyndman et al. (2008) introduced the Hurdle Shifted Poisson filter, aligning with Croston’s forecasts while accounting for count data. Building on this, Snyder et al. (2012) and Taylor (2012) proposed Poisson- and Negative Binomial-based filters, with the latter offering improved forecast accuracy. Moving beyond integer demand assumptions, Teunter et al. (2011) developed a dual SES approach forecasting demand size and occurrence probability directly, which has been used since then in a wider variety of contexts (for example, in Babai et al., 2014; Segerstedt and Leven, 2023; Doszyń and Dudek, 2024). Most recently, Svetunkov and Boylan (2023) introduced an ETS-based model combining Bernoulli and positive distributions (e.g., Gamma), demonstrating superior forecasting performance for both fractional and count data.
@@ 2.3.2 Combination approaches
Another key research stream in intermittent demand focuses on aggregation and combination approaches. Nikolopoulos et al. (2011) proposed aggregating intermittent demand to a regular demand level, forecasting it, and then disaggregating back to the original level, allowing the use of conventional forecasting methods without the common challenges of intermittent data. Petropoulos and Kourentzes (2015) extended this idea by combining forecasts across different aggregation levels. Kourentzes and Athanasopoulos (2021) applied temporal hierarchies to combine point forecasts, capturing hidden structures like trends and seasonality in intermittent demand. Finally, Wang et al. (2024) introduced probabilistic forecast combinations, showing that simple average combination performs best for quantile forecasts and inventory metrics.
@@ 2.3.3 Machine learning for intermittent demand
The paper by Kourentzes (2013) was the first one that we are aware of that used Artificial Neural Networks (ANN) for intermittent demand forecasting. He proposed two architectures: one capturing demand sizes and intervals in Croston’s style before dividing them, and another directly forecasting final demand. Both used demand sizes and intervals as inputs. While these methods did not outperform simpler forecasting approaches (e.g., Croston’s method), they showed slight improvements in inventory metrics.
@@
Nikolopoulos et al. (2016) evaluated performance of the k-nearest neighbour method for intermittent demand forecasting, comparing it with the conventional forecasting methods, showing that it outperforms Croston, SBA and SES on real data.
@@
Babai et al. (2020) compared SES, Croston, SBA, and bootstrapping methods with a Multilayer Perceptron ANN on a spare parts dataset, showing that a well-designed neural network could outperform conventional methods in point forecasts and inventory costs. Similarly, Türkmen et al. (2021) used a recurrent neural network with Negative Binomial and Poisson distributions for demand sizes and parametric distributions for demand intervals. Their approach, tested on several intermittent demand datasets, proved competitive with statistical methods, sometimes outperforming them in point forecasts and specific quantiles.
@@
Jiang et al. (2021) introduced an adaptive Support Vector Machine for spare parts demand forecasting, comparing it with parametric, bootstrap, and neural network methods. Their approach performed well in Mean Absolute Error (MAE) and scaled Mean Error, but the chosen error measures, minimised by the median, which in intermittent demand can often lead toward models predicting values closer to zero.
@@
Rožanec et al. (2022) proposed separating demand into regular and intermittent categories, using a gradient-boosted decision tree (CatBoost) to predict demand occurrence and a light gradient boosting machine (LightGBM) for demand sizes. Their approach outperformed conventional methods (e.g., Naïve, SES, SMA) in Area Under the Curve (AUC) and Mean Absolute Scaled Error (MASE). However, their methodology had drawbacks: (1) MASE is minimised by median, potentially selecting models biased toward zero demand; (2) instead of using probability of occurrence, the authors applied an arbitrary threshold to classify forecasts as zero or one. While suitable for classification, this is problematic for intermittent demand, because the occurrence of intermittent demand is fundamentally unpredictable. By forcing a binary classification, the model risked capturing noise rather than underlying patterns. Nonetheless, the idea of distinguishing between regular and intermittent demand remains valuable.
@@
Shrivastava et al. (2023) developed few Recurrent Neural Networks that would produce forecasts for demand sizes and probability of occurrence and demonstrated that this approach outperforms Croston, SBA and TSB on the M5 and car parts datasets.

### v2

@@ Abstract
Effective demand forecasting is critical for inventory management, production planning, and decision making across industries. Selecting the appropriate model and suitable features to efficiently capture patterns in the data is one of the main challenges in demand forecasting. In reality, this becomes even more complicated when the recorded sales have zeroes, which can happen naturally or due to some anomalies, such as stockouts and recording errors. Mistreating the zeroes can lead to the application of inappropriate forecasting methods, and thus leading to poor decision making. Furthermore, the demand itself can have different fundamental characteristics, and being able to distinguish one type from another might bring substantial benefits in terms of accuracy and thus decision making. We propose a two-stage model-based classification framework that in the first step, identifies artificially occurring zeroes, and in the second, classifies demand to one of the possible types: regular/intermittent, intermittent smooth/lumpy, fractional/count. The framework relies on statistical modelling and information criteria. We argue that different types of demand need different features, and show empirically that they tend to increase the accuracy of the forecasting methods and reduce inventory costs compared to those applied directly to the dataset without the generated features and the two-stage framework.
@@ 1 Introduction
People working in the area of demand forecasting sometimes encounter zeroes in their data. These zeroes can happen for a variety of reasons: some of them occur naturally, because no one buys the product in that specific moment of time, some happen artificially due to problems in inventory management system or recording errors. It is important to distinguish these two situations and treat them differently. In case of the naturally occurring zeroes, we typically have intermittent demand, which according to the definition of Svetunkov2023b is the demand that “has non-zero values occurring at irregular frequency”. Many statistical methods have been developed, starting from the Croston1972 and its bias corrected form by Syntetos2001 to machine learning methods developed over the years (Hasni2018; Babai2020; Jiang2021a), so an analyst can select the most appropriate or the favourite approach and use it for intermittent demand forecasting. However, if we deal with the artificially occurring zeroes, they need to be treated differently: for example, using an intermittent demand approach on the data with stockouts would be harmful for decision making, because we would be forecasting stockouts instead of demand.
@@
    suggesting several fundamental features based on (1) and (2) that, as we argue, should improve the performance of approaches in terms of forecasting and inventory management.
@@ 2.1 What is “intermittent demand”?
The rise of the interest in the area of intermittent demand started with the paper of Croston1972, who acknowledged that the simple exponential smoothing (SES by Brown1956) is not appropriate when used on the data with unpredictable zeroes. To solve the problem, he suggested to split the data into two time series: demand sizes and demand intervals. His idea was that the demand that we observe ($y_{t}$) can be represented as a combination of two variables:
@@
where $z_{t}$ is the demand size on observation $t$ and $o_{t}$ is a binary variable of demand occurrence, that has a probability of occurrence $p_{t}=\frac{1}{q_{t}}$, where $q_{t}$ is the interval between the observed demand sizes. To produce forecasts, Croston1972 used two simple exponential smoothing methods for capturing dynamics of the demand sizes and demand intervals. While being efficient and innovative, his approach was neglected by academia for more than 20 years until Willemain1994 and Johnston1996 showed that Croston’s method performed well in practice and should be preferred for intermittent demand instead of other simple forecasting methods. Acknowledging the existence of intermittent demand, these papers also opened a new direction of research – intermittent demand forecasting, where the patterns of the data are so different that the conventional forecasting techniques might fail or not work efficiently.
@@
From the practical point of view, when separating the regular demand from the intermittent, one still faces a challenge, because there are no appropriate rules and it is not clear, what quantity of zeroes transforms the regular demand into intermittent. Some practitioners use arbitrary thresholds of 10%, 15%, 20% etc of zeroes in the data as cut-off points, where one would need to switch from regular to intermittent demand forecasting method. But those thresholds do not have any theoretical rationale behind them, and can only be considered as approximations to the real solution of the problem. Even non-OR methods that could be used to forecast special events are still based on some arbitrary threshold values which differ across domains (Nikolopoulos2021). So far, the assumed “rule of thumb” is that any quantity of zeroes implies that the demand is intermittent.
@@
Furthermore, the academic definitions of intermittent demand found in the literature are generally broad. Early definitions, such as Silver1981, describe it as demand with “significant periods of no demand activity” but the term “significant” is unclear. Later definitions, including Willemain1994 and Syntetos2001, describe it as random demand with many (or large proportion of) zero periods, but the terms “many” and “large proportion” used remain ambiguous. More recent definitions of Syntetos2005, Syntetos2009, Teunter2011, and Babai2014a describe intermittent demand as appearing sporadically, which is, in essence, correct, but does not distinguish between random absence of demand and systematic gaps caused by external factors such as disruptions or recording errors.
@@
Lately, some definitions became more case-specific and sometimes confusing, like “the features of intermittent demand are characterised by their irregularity, with a very small demand size” (Jiang2021a), or that it is “characterised by time series with many zeroes” (Prestwich2021), or characterised by “irregular demand occurrence and low demand quantity variation” (Rozanec2022).
@@
One other mistake sometimes made in the literature is equating count demand to the intermittent one (Snyder2012). While in some situations this is correct, this is not universally the case because the term “intermittent” relates to the demand intervals, while the word “count” describes demand sizes. In reality, there are many examples of demand being intermittent yet fractional (e.g. electricity vehicles charging).
@@
Among all definitions, we find the one by Svetunkov2023b to be the most concise and clear: “intermittent time series is a series that has non-zero values occurring at irregular frequency”. This definition shows that the non-zero values happen randomly and cannot be predicted, and the definition does not impose any subjective terms like “many”, or “large”, or “some”. However, it still does not make a distinction between naturally occurring zeroes and the ones happening due to external factors.
@@ 2.2 Demand classification
There are several proposed classification schemes based either on variance partition, or the accuracy of forecasting procedures (Babiloni2010). We discuss the main ones here.
Williams1984 was one of the first who proposed classifying demand into ‘smooth’, ‘slow-moving’ and ‘sporadic’ by partitioning the variance of demand over the lead time into causal parts. Eaves2004 expanded this classification by inclusion of the irregular type, which is differentiated from the smooth one according to the level of demand sizes variability. However, both papers clearly state that the boundaries between these categories are industry/data-specific, and they should come as a managerial decision.
Syntetos2005a proposed a classification (now called ‘SBC’, Syntetos-Boylan-Croston) based on other principles. They discussed the existing categorisation schemes and compared Croston, Simple Exponential Smoothing (SES) and Syntetos-Boylan Approximation (SBA from Syntetos2005) based on their theoretical Mean Square Error (MSE) values. The authors showed that one can derive the cut-off values for average demand intervals and coefficient of variation, allowing to select between Croston and SBA for each type of demand. This scheme has gained large popularity among practitioners, because it suggest four distinct categories of intermittent demand:
@@
In many cases practitioners use the scheme as a prior step for data analysis, categorising the intermittent demand into these categories without any specific purpose, completely neglecting it in the following steps, where, for example, they use some machine learning methods for demand forecasting. We should point out that the original motivation of Syntetos2005a was to help in choosing between Croston and SBA, rather than doing a categorisation for the sake of it. This idea seems to have been lost over the years.
The modification of the SBC was proposed by Kostenko2006, who modified it, using an inequality from the ratio of MSEs from the original paper of Syntetos2005a. They demonstrated that the parameter space for different forecasting methods has a non-linear cut-off. This classification, while being mathematically correct, has not gained as much popularity as SBC, being more complex.
Finally, Petropoulos2015 proposed a further refinement of the Kostenko2006 scheme by adding SES method to the classification. They showed on the example of the Royal Airforce data (RAF) that this addition produces more accurate and less biased forecasts than the Kostenko2006 scheme.
However, all the research that has been done in this direction up to this point has mainly focused on selecting between Croston and SBA specifically for intermittent demand. This implies that if there is at least one zero in the data, then the data can be flagged as intermittent and either of these two methods should be used. While this is widely true, there is evidence from Syntetos2006 that SES and Simple Moving Average (SMA) perform well even on data with some zeroes. So a refined approach is needed in order to classify demand as intermittent or regular.
@@ 2.3.1 Exponential smoothing
Research on exponential smoothing for intermittent demand has advanced from early explorations of Croston’s method and its statistical foundations (Shenstone2005) to more refined models tailored to count and intermittent data. Hyndman2008b introduced the Hurdle Shifted Poisson filter, aligning with Croston’s forecasts while accounting for count data. Building on this, Snyder2012 and Taylor2012 proposed Poisson- and Negative Binomial-based filters, with the latter offering improved forecast accuracy. Moving beyond integer demand assumptions, Teunter2011 developed a dual SES approach forecasting demand size and occurrence probability directly, which has been used since then in a wider variety of contexts (for example, in Babai2014a; Segerstedt2023; Doszyn2024). Most recently, Svetunkov2023b introduced an ETS-based model combining Bernoulli and positive distributions (e.g., Gamma), demonstrating superior forecasting performance for both fractional and count data.
@@ 2.3.2 Combination approaches
Another key research stream in intermittent demand focuses on aggregation and combination approaches. Nikolopoulos2011a proposed aggregating intermittent demand to a regular demand level, forecasting it, and then disaggregating back to the original level, allowing the use of conventional forecasting methods without the common challenges of intermittent data. Petropoulos2015 extended this idea by combining forecasts across different aggregation levels. Kourentzes2021 applied temporal hierarchies to combine point forecasts, capturing hidden structures like trends and seasonality in intermittent demand. Finally, Wang2024 introduced probabilistic forecast combinations, showing that simple average combination performs best for quantile forecasts and inventory metrics.
@@ 2.3.3 Machine learning for intermittent demand
The paper by Kourentzes2013 was the first one that we are aware of that used Artificial Neural Networks (ANN) for intermittent demand forecasting. He proposed two architectures: one capturing demand sizes and intervals in Croston’s style before dividing them, and another directly forecasting final demand. Both used demand sizes and intervals as inputs. While these methods did not outperform simpler forecasting approaches (e.g., Croston’s method), they showed slight improvements in inventory metrics.
@@
Nikolopoulos2016 evaluated performance of the k-nearest neighbour method for intermittent demand forecasting, comparing it with the conventional forecasting methods, showing that it outperforms Croston, SBA and SES on real data.
@@
Babai2020 compared SES, Croston, SBA, and bootstrapping methods with a Multilayer Perceptron ANN on a spare parts dataset, showing that a well-designed neural network could outperform conventional methods in point forecasts and inventory costs. Similarly, Turkmen2019 used a recurrent neural network with Negative Binomial and Poisson distributions for demand sizes and parametric distributions for demand intervals. Their approach, tested on several intermittent demand datasets, proved competitive with statistical methods, sometimes outperforming them in point forecasts and specific quantiles.
@@
Jiang2021a introduced an adaptive Support Vector Machine for spare parts demand forecasting, comparing it with parametric, bootstrap, and neural network methods. Their approach performed well in Mean Absolute Error (MAE) and scaled Mean Error, but the chosen error measures, minimised by the median, which in intermittent demand can often lead toward models predicting values closer to zero.
@@
Rozanec2022 proposed separating demand into regular and intermittent categories, using a gradient-boosted decision tree (CatBoost) to predict demand occurrence and a light gradient boosting machine (LightGBM) for demand sizes. Their approach outperformed conventional methods (e.g., Naïve, SES, SMA) in Area Under the Curve (AUC) and Mean Absolute Scaled Error (MASE). However, their methodology had drawbacks: (1) MASE is minimised by median, potentially selecting models biased toward zero demand; (2) instead of using probability of occurrence, the authors applied an arbitrary threshold to classify forecasts as zero or one. While suitable for classification, this is problematic for intermittent demand, because the occurrence of intermittent demand is fundamentally unpredictable. By forcing a binary classification, the model risked capturing noise rather than underlying patterns. Nonetheless, the idea of distinguishing between regular and intermittent demand remains valuable.
@@
Shrivastava2023 developed few Recurrent Neural Networks that would produce forecasts for demand sizes and probability of occurrence and demonstrated that this approach outperforms Croston, SBA and TSB on the M5 and car parts datasets.
In a recent paper Zhang2024 use Transformer architecture to forecast intermittent demand. While the idea is interesting, the main flaw of the paper is in the evaluation, where the authors violate the basic forecast evaluation principles (no explanation of the forecast horizon, inadequate error measures used etc).